# How simple can you go? An off-the-shelf transformer approach to molecular dynamics

### v2

@@ Acknowledgements
We thank Luis Müller for his encouragement and being a great colleague. We also thank Elron Pens, Stefan Chmiela and Sidney Bender for important insights and fruitful discussions. Stefan Gugler was supported by the Postdoc.Mobility fellowship by the Swiss National Science Foundation (project no. 225476)
@@ Appendix A Appendix
A.1 Side-by-side comparison
Table 5: Comparison of different force field approaches: MD-ET, MLFF (typical learned force fields), and Traditional FF (classical potentials).
Feature MD-ET MLFF Traditional FF
Equivariance (close to numerical precision)
Energy Conservation
Short NVE Stability $\sim$
Long NVE Stability
Short NVT Stability
Long NVT Stability $\sim$
Differentiable
Fast (per step)
Accurate (quantum-level)
Low Inductive Bias $\sim$
A.2 Edge Transformer Implementation
A.2.1 Molecular embedding layer details
To create initial edge representations, we combine several embeddings: i) spin & charge, ii) atomic numbers , iii) pairwise distances, and iv) pairwise directions.
@@ A.2.1 Molecular embedding layer details
$\mathbf{z}_{i}=\mathbf{W}^{\text{spin}}{h}(s_{i})+\mathbf{W}^{\text{charge}}{h}(q_{i}),$ (7)
where $\mathbf{W}^{\text{spin}}\in\mathbb{R}^{D\times d_{\text{spin}}}$ and $\mathbf{W}^{\text{charge}}\in\mathbb{R}^{D\times d_{\text{charge}}}$ are learned weight matrices for spin and charge, respectively, ${h}$ is a one-hot encoding and $d_{\text{spin}}$ and $d_{\text{charge}}$ are the sizes of the spin and charge vocabularies, respectively.
@@
$A_{ij}=\psi_{\text{edge}}\Big{(}$ $\left[\mathbf{W}^{\text{atom}}{h}(Z_{i})+\mathbf{W}^{\text{cfg}}\tau(Z_{i})+\mathbf{z}_{i},\right.$
$\left.\mathbf{W}^{\text{atom}}{h}(Z_{j})+\mathbf{W}^{\text{cfg}}\tau(Z_{j})+\mathbf{z}_{j}\right]\Big{)}\ ,$ (8)
where $Z_{i},Z_{j}\in\mathbb{N}$ are the atomic numbers and $\mathbf{W}^{\text{atom}}\in\mathbb{R}^{D\times d_{\text{atom}}}$ and $\mathbf{W}^{\text{cfg}}\in\mathbb{R}^{D\times d_{\text{cfg}}}$ are the learned weight matrices for atomic numbers and electron configuration, $\tau$ (see Appendix A.2.1), respectively. Lastly, $\psi_{\text{edge}}:\mathbb{R}^{2D}\rightarrow\mathbb{R}^{D}$ is an MLP that projects concatenated embeddings into the embedding space.
@@
where $\psi_{\text{dir}}$ is another learnable linear layer.
@@
Prediction Head & Training Objective
@@
Electronic embedding
The function $\tau:\mathbb{N}\rightarrow\mathbb{R}^{d_{\text{cfg}}}$ maps an atomic number $Z_{i}$ to a fixed electron configuration descriptor that encodes i) subshell occupancy, the number of electrons in each $s$, $p$, $d$, and $f$ subshell up to the valence shell, ii) valence electrons in the outermost $s$ (vs), $p$ (vp), $d$ (vd), and $f$ (vf) subshells.
This descriptor captures an atom’s electronic structure, which governs its chemical behavior. For example, chlorine ($Z=17$) and bromine ($Z=35$) both have $7$ valence electrons ($\text{vs}=2$, $\text{vp}=5$), explaining their tendency to form -1 ions. The configuration follows the Aufbau principle, with exceptions in transition metals and lanthanides/actinides accounted for empirically.
$\tau(Z_{i})$ retrieves a precomputed vector from a lookup table (excerpt below). Each row corresponds to an element’s ground-state electron configuration and valence electrons. The full table spans $Z=1$ (H) to $Z=100$ (Fm).
Table 6: Electron configuration descriptor examples. Full table available in source code.
Z Element 1s 2s 2p 3s 3p vs/vp/vd/vf
1 H 1 0 0 0 0 1/0/0/0
8 O 2 2 4 0 0 2/4/0/0
17 Cl 2 2 6 2 5 2/5/0/0
35 Br 2 2 6 2 6 2/5/10/0
79 Au 2 2 6 2 6 1/0/10/14
The descriptor vector is linearly transformed by $\mathbf{W}^{\text{cfg}}\in\mathbb{R}^{D\times d_{\text{cfg}}}$ to produce a $D$-dimensional embedding. This allows the model to learn similarities between elements with analogous configurations (e.g., halogens, noble gases) directly from their electronic structure.
@@ 79 Au 2 2 6 2 6 1/0/10/14
We implement MD-ET using PyTorch 2.5.1. Pseudo-code of the triangular and a comparison with regular (single-head) attention can be found in Algorithm 1 (below). To improve performance, we compile the triangular attention operation at runtime using torch compile. This necessitates rewriting triangular attention without einsums.
A.3 Inference Speed Evaluation
See Table 7 below. For runtime benchmarks, we time the forward pass of the model, including all postprocessing steps on a cloud-based V100 GPU (Throughput, Latency). Inference Time measures the pure inference time without postprocessing. To give an approximation of the performance impact of frame-averaging we report the model throughput for multiple batch sizes and structures of different sizes. To make the results more comparable to different hardware we also provide total FLOPs per forward pass.
$\begin{array}[]{ccccccc}\hline\cr\hline\cr\text{Molecule}&\text{Batch Size}&\text{Throughput}&\text{Avg Latency (ms)}&\text{Inference Time (ms)}&\text{FLOPs}\\ \hline\cr\text{Aspirin}&1&159.800&6.258&4.927&5.42\times 10^{9}\\ &2&157.909&6.333&5.063&1.08\times 10^{10}\\ &4&147.431&6.783&5.505&2.17\times 10^{10}\\ &8&102.680&9.739&8.458&4.33\times 10^{10}\\ &16&63.636&15.714&14.414&8.67\times 10^{10}\\ &32&35.745&27.976&26.652&1.73\times 10^{11}\\ &64&18.685&53.520&52.103&3.47\times 10^{11}\\ \hline\cr\text{Ethanol}&1&148.151&6.750&5.429&9.96\times 10^{8}\\ &2&146.148&6.842&5.586&1.99\times 10^{9}\\ &4&151.253&6.611&5.343&3.98\times 10^{9}\\ &8&153.357&6.521&5.247&7.97\times 10^{9}\\ &16&151.867&6.585&5.313&1.59\times 10^{10}\\ &32&115.476&8.660&7.377&3.19\times 10^{10}\\ &64&70.741&14.136&12.784&6.37\times 10^{10}\\ \hline\cr\text{Naphthalene}&1&150.254&6.655&5.320&3.98\times 10^{9}\\ &2&151.231&6.612&5.333&7.96\times 10^{9}\\ &4&151.409&6.605&5.332&1.59\times 10^{10}\\ &8&114.743&8.715&7.458&3.18\times 10^{10}\\ &16&70.293&14.226&12.937&6.37\times 10^{10}\\ &32&41.122&24.318&22.987&1.27\times 10^{11}\\ &64&22.033&45.387&43.976&2.55\times 10^{11}\\ \hline\cr\text{Salicylic Acid}&1&147.294&6.789&5.481&3.14\times 10^{9}\\ &2&150.733&6.634&5.371&6.29\times 10^{9}\\ &4&154.547&6.471&5.225&1.26\times 10^{10}\\ &8&130.698&7.651&6.407&2.52\times 10^{10}\\ &16&85.649&11.676&10.401&5.03\times 10^{10}\\ &32&51.222&19.523&18.236&1.01\times 10^{11}\\ &64&27.834&35.927&34.564&2.01\times 10^{11}\\ \hline\cr\hline\cr\end{array}$
@@ A.3 Inference Speed Evaluation
A.4 Equivariance proof
See main text for the equivariance error, Eq. (5).
Proof: We first simplify the first term and substitute the equivariance property $f_{\theta}(\mathcal{R}\boldsymbol{x})=\mathcal{R}f_{\theta}(\boldsymbol{x})$:
$\mathbb{E}_{\mathcal{R}}\left[\mathcal{R}^{\top}f_{\theta}(\mathcal{R}\boldsymbol{x})\right]=\mathbb{E}_{\mathcal{R}}\left[\mathcal{R}^{\top}\mathcal{R}f_{\theta}(\boldsymbol{x})\right].$ (17)
Since $\mathcal{R}^{\top}\mathcal{R}=\boldsymbol{I}$ (identity matrix) for all $\mathcal{R}\in\text{SO}(3)$:
$\mathbb{E}_{\mathcal{R}}\left[\boldsymbol{I}f_{\theta}(\boldsymbol{x})\right]=\mathbb{E}_{\mathcal{R}}\left[f_{\theta}(\boldsymbol{x})\right].$ (18)
The expectation $\mathbb{E}_{\mathcal{R}}[f_{\theta}(\boldsymbol{x})]$ is over $\mathcal{R}$, but $f_{\theta}(\boldsymbol{x})$ does not depend on $\mathcal{R}$. So we obtain
For the 2nd term we apply the equivariance property to $f_{\theta}(\mathcal{S}\boldsymbol{x})$
$f_{\theta}(\mathcal{S}\boldsymbol{x})=\mathcal{S}f_{\theta}(\boldsymbol{x}).$ (20)
and similarly obtain
Therefore, the entire expectation over $x$ and $\mathcal{S}$ is
$\mathbb{E}_{x,\mathcal{S}}[||f_{\theta}(\boldsymbol{x})-f_{\theta}(\boldsymbol{x})||]=0.$ (22)
2nd Claim: $\lim_{n\to\infty}E_{\text{eq}}(\mathcal{D},\hat{f}^{\prime})=0$ for all neural nets.
Proof: We insert $\hat{f_{\theta}}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{R}_{i}^{\top}\hat{f}_{\theta}(\mathcal{R}_{i}x)$ into $\mathcal{S}^{\top}f_{\theta}(\mathcal{S}\boldsymbol{x})$, obtaining:
$\mathcal{S}^{\top}\frac{1}{n}\sum_{i=1}^{n}\mathcal{R}_{i}^{\top}\hat{f}_{\theta}(\mathcal{R}_{i}\mathcal{S}\boldsymbol{x})$ (23)
Rearranging, we obtain:
$\frac{1}{n}\sum_{i=1}^{n}(\mathcal{R}_{i}\mathcal{S})^{\top}\hat{f}_{\theta}(\mathcal{R}_{i}\mathcal{S}\boldsymbol{x})$ (24)
We now substitute $\mathcal{R}_{i}\mathcal{S}=\mathcal{T}_{i}\in SO(3)$, obtaining
$\frac{1}{n}\sum_{i=1}^{n}(\mathcal{T}_{i})^{\top}\hat{f}_{\theta}(\mathcal{T}_{i}\boldsymbol{x})$ (25)
According to Kolmogorov’s law the sample mean converges to the expectation for large numbers of samples. For $\lim_{n\rightarrow\infty}$ We therefore obtain
$\mathbb{E}_{\mathcal{T}\sim SO(3)}[\mathcal{T}^{\top}\hat{f}_{\theta}(\mathcal{T}\boldsymbol{x})]$ (26)
Inserting into equation 5 it is trivial to see that $E_{\text{eq}}(\mathcal{D},\hat{f}^{\prime})=0$.
@@ A.6 Experimental Setup for Benchmarks
All fine-tuning experiments employ the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay $\lambda=1e-7$. Experiments were conducted on a single NVIDIA H100 GPU with a batch size of 500. When the full batch exceeded GPU VRAM, gradient accumulation was used. The learning rate schedule combined cosine annealing with a linear warmup: fine-tuning lasted 2000 steps if not specified differently, with the first 100 steps as warmup, and a minimum learning rate of 1e-8. During fine-tuning, each sample was augmented with two transformations—random rotations and random reflections.
A.6.1 MD17
For finetuning on the MD17 experiments we follow the experimental setup of [58], namely we sample 9500/500/10000 conformations for train/val/test splits for each molecule and evaluate the force prediction error on the test set. The hyperparameters for each molecule are shown in table 9.
@@ A.6.1 MD17
For the stability evaluation we run an MD simulation with a timestep of 0.5fs for 600k steps. We use a Nosé-Hoover thermostat set to 500K. To improve predictions, we use frame-averaging with $n=4$ samples per step. For more details, see [58].
@@ A.6.3 Ko2020
A.7 Experimental setup for evaluation of energy conservation
For the evaluation of energy conservation we run NVE simulations for 1 million using a step size of 0.5fs. To improve predictions, we use frame-averaging with $n=16$ samples per step.
Appendix B Visualization of molecules
Figure 4: Visualizations of the molecular structures used in this work. In the upper half are the cumulenes (left) and the alkanes (right) with increasing carbon chain length. On the bottom left are the four molecules used from MD17: naphthalene, aspirin, salicylic acid, and ethanol. On the bottom right are the four systems studied by Ko et al. [53]: a gold dimer on a magnesium oxide surface, a silver trimer, a salt system, and $\text{C}_{10}\text{H}_{2}$ (corresponding $\text{C}_{10}\text{H}^{+}$ omitted).
Algorithm 1 Comparison between standard attention and triangular attention
  function ATTENTION($X:n\times d$)
   $Q,K,V\leftarrow\text{linear}(X).\text{chunk}(3)$
   $\tilde{A}\leftarrow\text{einsum}(id,jd\rightarrow ij,Q,K)$
   $A\leftarrow\text{softmax}(\tilde{A}/\sqrt{d},-1)$
   $O\leftarrow\text{einsum}(ij,jd\rightarrow id,A,V)$
   return $\text{linear}(O)$
  end function
  function TRI_ATTENTION($X:n\times n\times d$)
   $Q,K,V^{1},V^{2}\leftarrow\text{linear}(X).\text{chunk}(4)$
   $V\leftarrow\text{einsum}(ild,ljd\rightarrow iljd,V^{1},V^{2})$
   $\tilde{A}\leftarrow\text{einsum}(ild,ljd\rightarrow ilj,Q,K)$
   $A\leftarrow\text{softmax}(\tilde{A}/\sqrt{d},-1)$
   $O\leftarrow\text{einsum}(ij,iljd\rightarrow ijd,A,V)$
   return $\text{linear}(O)$
  end function

### v3

@@ Acknowledgements
We thank Luis Müller for his encouragement and being a great colleague. We also thank Elron Pens, Stefan Chmiela, and Sidney Bender for important insights and fruitful discussions. SGu was supported by the Postdoc.Mobility fellowship of the Swiss National Science Foundation (project no. 225476). ME, SGu, and KRM acknowledge support by the German Ministry of Education and Research (BMBF) for BIFOLD (01IS18037A). Further, this work was in part supported by the BMBF under Grants 01IS14013A-E, 01GQ1115, 01GQ0850, 01IS18025A, and 031L0207D. KRM was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korea government (MSIT) (No.2019- 0-00079, Artificial Intelligence Graduate School Program, Korea University and No. 2022-0-00984, Development of Artificial Intelligence Technology for Personalized Plug-and-Play Explanation and Verification of Explanation).
@@ Appendix A Appendix
A.1 Supplementary Figures
Figure 8: Total energy variation per atom over time for NVE simulations of additional alkanes, using MD-ET evaluated with fp64 precision.
Figure 9: Development of equivariance error $E_{\text{eq}}$ and potential energy along a 500 ps NVE trajectory simulating $\text{C}_{8}\text{H}_{18}$ sampled every 50 fs. Comparison between MD-ET (identical checkpoint, fp32 and fp64 evaluated) and Orb-v2. The Orb plot only shows the first 50 ps since the molecule “explodes” after, which makes energy calculations unreliable. Potential energy of frames calculated with PBE def2-SVP.
Figure 10: Velocity-velocity autocorrelation power spectra of a 1 ns simulation of $\text{C}_{8}\text{H}_{18}$ using the SVR thermostat with $\tau=1000$. Comparison of Orb and MD-ET (fp64 evaluated) with an experimentally observed infrared absorption spectrum [8, 7]
Figure 11: Mean antisymmetric ratio $\lambda$ of force predictions of 2048 QCML train/test samples. Sizes smaller than 2 atoms or with less than 10 samples have been omitted.
A.2 Molecular embedding layer implementation details
To create initial edge representations, we combine several embeddings: i) spin and charge, ii) atomic numbers, iii) pairwise distances, and iv) pairwise directions.
@@ A.2 Molecular embedding layer implementation details
$\mathbf{e}_{i}=\mathbf{W}^{\text{spin}}{h}(s_{i})+\mathbf{W}^{\text{charge}}{h}(q_{i}),$ (7)
where $\mathbf{W}^{\text{spin}}\in\mathbb{R}^{D\times d_{\text{spin}}}$ and $\mathbf{W}^{\text{charge}}\in\mathbb{R}^{D\times d_{\text{charge}}}$ are learned weight matrices for spin and charge, respectively, h{h} is a one-hot encoding and $d_{\text{spin}}$ and $d_{\text{charge}}$ are the sizes of the spin and charge vocabularies, i.e., how many distinct spin states and charge states are possible, respectively.
@@
$A_{ij}=\psi_{\text{edge}}\Big($ $\left[\mathbf{W}^{\text{atom}}{h}(Z_{i})+\mathbf{e}_{i},\right$.
$\left.\mathbf{W}^{\text{atom}}{h}(Z_{j})+\mathbf{e}_{j}\right]\Big)\ ,$ (8)
where $Z_{i},Z_{j}\in\mathbb{N}$ are the atomic numbers and $\mathbf{W}^{\text{atom}}\in\mathbb{R}^{D\times d_{\text{atom}}}$ are the learned embedding weights for the atomic numbers. Lastly, $\psi_{\text{edge}}:\mathbb{R}^{2D}\rightarrow\mathbb{R}^{D}$ is an MLP that projects concatenated embeddings into the embedding space.
@@
where $\psi_{\text{dir}}$ is another linear layer.
@@
A.3 MD-ET implementation details
Prediction head and training objective
@@ A.3 MD-ET implementation details
We implement MD-ET using PyTorch 2.5.1. Pseudo-code of the triangular attention mechanism and a comparison with regular (single-head) attention can be found in Algorithm 1 (below). To improve performance, we compile the triangular attention operation at runtime using torch.compile. This necessitates rewriting triangular attention without Einstein summation, einsum.
Algorithm 1 Comparison of standard attention vs. triangular attention (TRIA)
1:function ATTENTION($\boldsymbol{X}:N\times D$)
2:  $\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V}\leftarrow\text{linear}(\boldsymbol{X}).\text{chunk}(3)$
3:
4:  $\tilde{\boldsymbol{A}}\leftarrow\text{einsum}(id,jd\rightarrow ij,\boldsymbol{Q},\boldsymbol{K})$
5:  $\boldsymbol{A}\leftarrow\text{softmax}\left(\frac{\tilde{\boldsymbol{A}}}{\sqrt{d}}\right)$
6:  $\boldsymbol{O}\leftarrow\text{einsum}(ij,jd\rightarrow id,\boldsymbol{A},\boldsymbol{V})$
7:  return $\text{linear}(\boldsymbol{O})$
8:end function
9:function TRIA($\boldsymbol{X}:N\times N\times D$) 10:  $\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V}^{1},\boldsymbol{V}^{2}\leftarrow\text{linear}(\boldsymbol{X}).\text{chunk}(4)$ 11:  $\boldsymbol{V}\leftarrow\boldsymbol{V}^{1}_{il}\odot\boldsymbol{V}^{2}_{lj}$ # Element-wise product 12:  $\tilde{\boldsymbol{A}}\leftarrow\text{einsum}(ild,ljd\rightarrow ilj,\boldsymbol{Q},\boldsymbol{K})$ 13:  $\boldsymbol{A}\leftarrow\text{softmax}\left(\frac{\tilde{\boldsymbol{A}}}{\sqrt{d}}\right)$ 14:  $\boldsymbol{O}\leftarrow\text{einsum}(ij,iljd\rightarrow ijd,\boldsymbol{A},\boldsymbol{V})$ 15:  return $\text{linear}(\boldsymbol{O})$ 16:end function
A.4 Inference speed evaluation
Figure 12: Empirical inference speed and memory footprint scaling of MD-ET.
Figure 12 shows empirical measurements of inference time and memory footprint scaling on an H100 PCIE GPU. We can see that MD-ET’s cubic scaling becomes noticeable for system sizes larger than 100 atoms and problematic for systems larger than 500 atoms. See Table 7 below for runtime benchmarks on MD17. Here, we time the forward pass of the model, including all postprocessing steps on a cloud-based V100 GPU (following the benchmark described in [4]). Inference Time measures the pure inference time without postprocessing.
[…]
@@ A.6 Experimental setup for benchmarks
All fine-tuning experiments employ the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay $\lambda=1\times 10^{-7}$. Experiments were conducted on a single NVIDIA H100 GPU with a batch size of 500. When the full batch exceeded GPU VRAM, gradient accumulation was used. The learning rate schedule combined cosine annealing with a linear warmup: fine-tuning lasted 2000 steps if not specified differently, with the first 100 steps as warmup, and a minimum learning rate of $1\times 10^{-8}$. During fine-tuning, each sample was augmented with two transformations—random rotations and random reflections.
A.6.1 MD17-10k
For fine-tuning on the MD17 experiments we follow the experimental setup of [4], namely we sample 9500/500/10000 conformations for train/val/test splits for each molecule and evaluate the force prediction error on the test set. The hyperparameters for each molecule are shown in Table 9.
@@ A.6.1 MD17-10k
For the stability evaluation we run an MD simulation with a timestep of 0.5 fs for 600k steps. We use a Nosé–Hoover thermostat set to 500K. For more details, see [4].
@@ A.6.3 Ko2020
A.6.4 xxMD
For the experiments with the xxMD dataset we use the canonical splits of the xxMD-DFT subset provided by the authors. For fine-tuning experiments we use identical settings as for all other fine-tuning experiments, optimizing for 2000 steps. For training from scratch we use almost identical parameters to the QCML pretraining reported above, but reduce compute requirement by training for only 250,000 steps, using a batch size of 256 and a feed-forward layer multiplier of 2. We only hyper-optimize the learning rate, optimal learning rates are reported in table 12.
Table 12: Fine-tuning Hyperparameters for xxMD
Subset Learning Rate Direct Training Learning Rate fine-tuning
Azobenzene $5\times 10^{-4}$ $5\times 10^{-5}$
Dithiophene $5\times 10^{-4}$ $1\times 10^{-4}$
Malonaldehyde $5\times 10^{-4}$ $1\times 10^{-4}$
Stilbene $5\times 10^{-4}$ $1\times 10^{-4}$
A.6.5 SPICE
We use the SPICEv1 dataset with a random 90%/5%/5% split on the molecular level. Since our QCML-trained model is able to represent most elements up to Astatine, we do not filter based on molecular composition. We only filter conformers with an atomic force vector larger than 50 eV. See Table 13 for details on filtered samples per subset. In total we use around 98% of samples (compared to 85% for EScAIP and MACE-OFF[6]). We fine-tune for 2000 steps and use the same hyperparameters as for other datasets above, only tuning the learning rate (see Table 14).
Table 13: Filtered conformers from the SPICE dataset. Overall: 22900/1127651 samples filtered (2.03%)
Subset Train+Val Test
Filtered/Total % Filtered/Total %
Pubchem 21495/695212 3.09 1216/36644 3.32
Monomers 0/17700 0.00 0/900 0.00
Dimers 0/328020 0.00 0/14025 0.00
Dipeptides 187/32100 0.58 2/1750 0.11
Solvated Amino Acids 0/1200 0.00 0/100 0.00
Total 21682/1074232 2.02 1218/53419 2.28
Table 14: Fine-tuning Hyperparameters for SPICE
Subset Train Split Size Val Split Size Test Split Size Learning Rate
Pubchem 638,140 35,577 35,428 $5\times 10^{-4}$
Monomers 16,800 900 900 $1\times 10^{-4}$
Dimers 314,410 13,610 14,025 $5\times 10^{-4}$
Dipeptides 30,263 1,650 1,748 $5\times 10^{-4}$
Solvated Amino Acids 1,150 50 100 $1\times 10^{-4}$
A.7 Equivariance proof
See main text for the equivariance error, Eq. (5), which is 0 for an equivariant model. We first simplify the first term and substitute the equivariance property $f_{\theta}(\mathcal{R}\boldsymbol{x})=\mathcal{R}f_{\theta}(\boldsymbol{x})$:
$\mathbb{E}_{\mathcal{R}}\left[\mathcal{R}^{\top}f_{\theta}(\mathcal{R}\boldsymbol{x})\right]=\mathbb{E}_{\mathcal{R}}\left[\mathcal{R}^{\top}\mathcal{R}f_{\theta}(\boldsymbol{x})\right]=\mathbb{E}_{\mathcal{R}}\left[\boldsymbol{I}f_{\theta}(\boldsymbol{x})\right]=\mathbb{E}_{\mathcal{R}}\left[f_{\theta}(\boldsymbol{x})\right]$ (17)
since $\mathcal{R}^{\top}\mathcal{R}=\boldsymbol{I}$ (identity matrix) for all $\mathcal{R}\in\text{SO}(3)$. The expectation $\mathbb{E}_{\mathcal{R}}[f_{\theta}(\boldsymbol{x})]$ is over $\mathcal{R}$, but $f_{\theta}(\boldsymbol{x})$ does not depend on $\mathcal{R}$. So we obtain
For the 2nd term we also apply the equivariance property. We obtain
Therefore, the entire expectation over $x$ and $\mathcal{S}$ is
[…]
Appendix B Uniform sampling of rotations via the 600-cell
To obtain a set of 60 regularly spaced rotations in $\text{SO}(3)$ we use the 600-cell, a convex regular 4-dimensional polytope with 120 vertices. Since the 4 dimensional unit sphere $\mathbb{S}^{4}$ covers SO(3) exactly twice (double cover), $\mathbb{S}^{3}/{\pm 1}\cong\text{SO}(3)$, the quaternion representations $(w,x,y,z)$ and its antipode (w,x,y,z)-(w,x,y,z) map to the same 3D rotation. We can thus use the coordinates of half of the 600-cell’s 120 vertices to derive 60 uniformly spaced SO(3) rotations.
For the larger set of 360 uniform points, we combine the 60 points above with half of the 600-cell’s 600 tetrahedral faces (each defined by four vertices spanning a 3-simplex) and compute their centroids. Under the double cover, antipodal centroids map to the same $\text{SO}(3)$ rotation, yielding 300 points. The union of these two subsets forms a 360-point configuration in $\text{SO}(3)$ and is uniform due to the 600-cell’s symmetry.
Appendix C Comparison of 50 ns free energy surface on alanine dipeptide
To assess whether and how conformational distributions are impacted by the lack of energy conservation, we run a 50 ns MD simulation of alanine dipeptide (more precisely N-acetyl-L-alanine-N’-methylamide) both with the baseline MD-ET model as well as an identical model using automatic differentiation to obtain energy-conserving force predictions. To obtain this model, we first train an MD-ET architecture with both a force and an energy prediction head for 250k steps. Note that in order to obtain stable second-order gradients we need to replace the directional embedding procedure based on azimuth and polar angles with an MLP, directly embedding normalized displacement vectors. This has only a minor impact on performance. After 250k steps of direct training we remove the force prediction head and replace the energy prediction head with one deriving energy-conserving forces using automatic differentiation. We then fine-tune this modified model for an additional 50,000 steps as described in Bigi et al. [1] and [3].
We run 50 ns MD simulations with a step size of 0.5 fs using the SVR thermostat set to a target temperature of 500 K. We choose this temperature since the baseline model evaluated at fp32 precision shows a small but significant energy drift. To assess whether the random offset rotations described in Section 4.3 impact observables, we run an additional 50 ns simulation using the baseline model but with random offset rotations. The resulting free energy surfaces show no significant differences (see Figure 13).
Figure 13: Comparison of 50 ns free energy surfaces of alanine dipeptide at 500 K using the SVR thermostat. Left: Baseline MD-ET model; middle: Baseline MD-ET model using random rotational offsets in each frame; right: Identical network architecture trained on identical data using energy-conserving forces.